OpenStudy (anonymous):

R=log( I/Io) where Io is the intensity of a certain very small earthquake. What is the magnitude of an earthquake with intensity 8000 times Io?

OpenStudy (anonymous):

I=8000I0 R=log(II0) R=log(8000I0I0) R=log(8000) R=log(8)+log(103) R=3+log(8) R≈3.9

OpenStudy (anonymous):

Oh I got the log(103) when I divided

OpenStudy (anonymous):

just wondering... where exactly did the division occur.. did you replace the first I with the 8000? cause i dk what exactly the i0 is

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what do i replace the Io with? should i replace it with 1?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

oh ok, thank you!

OpenStudy (anonymous):

wait, when you do the math where exactly did you get the result 103.. cause i can't figure out what you divided to get it.. extra slow in math here =/

OpenStudy (anonymous):

It's hard to explain

OpenStudy (anonymous):

my math is different from the teachers but I still get the same answer

OpenStudy (anonymous):

i can see it is a bit different but in the line R=log(8000I0I0), do you divide log of 8000 * I o, and then divide it by I o or how does that change?

OpenStudy (anonymous):

right you are doing it

OpenStudy (anonymous):

so then the I o has to be a different number than 1 because 8000 *1 /1 = 8000 xD

OpenStudy (anonymous):

well played yes

OpenStudy (anonymous):

so what would that number be?

OpenStudy (anonymous):

Idk never did it that way

OpenStudy (anonymous):

so in that case how would you go about solving the problem?

OpenStudy (anonymous):

Like how I did up there

OpenStudy (anonymous):

what i'm asking is how did you get from R=log(8000I0I0) R=log(8000) to R=log(8)+log(103)

OpenStudy (anonymous):

division

OpenStudy (anonymous):

lol what exactly did you divide O.O

OpenStudy (anonymous):

R=log(8000I0I0) R=log(8000)

OpenStudy (anonymous):

see those I marks those are getting rid of the zeros making it 8000

OpenStudy (anonymous):

ooohhhh i see exactly what you mean! thanks man!

OpenStudy (anonymous):

your welcome :D